<html>
  <head>

    <meta http-equiv="content-type" content="text/html; charset=utf-8">
  </head>
  <body bgcolor="#FFFFFF" text="#000000">
    Folks,<br>
    <br>
    Please find the notes below and at<br>
    <br>
     
<a class="moz-txt-link-freetext" href="https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_October_6,_2014">https://halldweb1.jlab.org/wiki/index.php/Private:Hall_D_Staff_Meeting,_October_6,_2014</a><br>
    <br>
      -- Mark<br>
    _______________________________<br>
    <br>
    <div id="globalWrapper">
      <div id="column-content">
        <div id="content">
          <h1 id="firstHeading" class="firstHeading">Private:Hall D
            Staff Meeting, October 6, 2014</h1>
          <div id="bodyContent">
            <p>Present: Fernando Barbosa, Eugene Chudakov (chair), Mark
              Dalton, Alexandre Deur, Hovanes Egiyan, Sergey Furletov,
              Mark Ito, David Lawrence, Yi Qiang, Elton Smith, Sascha
              Somov, Simon Taylor, Beni Zihlmann<br>
            </p>
            <br>
            <span class="mw-headline" id="Eugene"></span>
            <h1><span class="mw-headline" id="Eugene">Eugene</span></h1>
            <ul>
              <li> Appraisals will be initiated today. They are due by
                Oct. 14.
                <ul>
                  <li> With uncertainty due to the upcoming continuing
                    resolution Physics had decided no promotions for now
                  </li>
                </ul>
              </li>
              <li> 12 GeV budget: project is done
                <ul>
                  <li> all construction accounts close
                  </li>
                  <li> one exception is account for pre-ops
                  </li>
                  <li> Eugene will issue guidance on how to use it
                  </li>
                  <li> should be used for paying non-Hall-D groups at
                    the lab
                  </li>
                </ul>
              </li>
              <li> Next Long-Range Plan meeting in April
              </li>
              <li> New PAC members: Frank Maas, Dean [?], Renee Fatemi,
                Werner Boeglin
              </li>
              <li> Post-doc search: choice sent to HR
              </li>
            </ul>
            <h1> <span class="mw-headline" id="David">David</span></h1>
            <ul>
              <li> work needed on translation tables
              </li>
              <li> auto-transfer to tape library needs directory
                handling logic included
              </li>
              <li> online monitoring: getting help from offline folks:
                Paul Mattione, Sean Dobbs, Kei Moriya
              </li>
              <li> need to modify read-out list to get DAQ parameters
                into data stream
              </li>
              <li> EPICS variables to data stream needs work
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Sergey_F.">Sergey F.</span></h1>
            <ul>
              <li> DAQ tested with CDC, FDC, BCAL, TOF, ST, FCAL in
                read-out
              </li>
              <li> read-out lists are scattered, need to consolidate
              </li>
              <li> highest rate so far 100 Hz with test pulser
                <ul>
                  <li> Sascha reports DAQ crashes on DMA from F125s
                  </li>
                </ul>
              </li>
              <li> monitoring of rates, occupancy being developed
              </li>
              <li> DAQ has gone at 17 kHz with no VME transfers
              </li>
              <li> Eugene asked about tagger hall readout
                <ul>
                  <li> Sascha: hodoscope has been read-out, worked with
                    Nathan Sparks
                  </li>
                  <li> S: microscope, jumper settings for addresses for
                    FADCs wrong, Sergey P. is fixing
                  </li>
                </ul>
              </li>
              <li> DAQ settings and pedestals can be set for all F250s
                and F125s, settings dumped to file
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Sascha">Sascha</span></h1>
            <ul>
              <li> working on global trigger for BCAL
                <ul>
                  <li> using LED to fire trigger
                  </li>
                  <li> CDC had been added
                  </li>
                  <li> failed at 400 Hz
                  </li>
                </ul>
              </li>
              <li> finalizing read-out list for all detectors
              </li>
              <li> want set-ups to be able to run with triggers and
                crates
              </li>
              <li> next plan: use same global energy for FCAL
                <ul>
                  <li> will first check in local mode
                  </li>
                  <li> timing issues in combining BCAL and FCAL
                  </li>
                </ul>
              </li>
              <li> next area of work: pair spectrometer
                <ul>
                  <li> trigger through GTP, course counters only
                  </li>
                </ul>
              </li>
              <li> Eugene: can specific crates be excluded from read-out
                based in trigger type?
                <ul>
                  <li> S: in principle yes, not implemented
                  </li>
                </ul>
              </li>
              <li> Vladimir Berdnikov has seen PS signals, looked at
                SiPM noise for each bias group
                <ul>
                  <li> read-out is done
                  </li>
                </ul>
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Fernando">Fernando</span></h1>
            <ul>
              <li> the Hall of Shame is expanding
                <ul>
                  <li> rack moved on upstream platform
                  </li>
                </ul>
              </li>
              <li> cables are all settled except for long-term network
                switch
                <ul>
                  <li> will need a PC to control microscope electronics
                  </li>
                </ul>
              </li>
              <li> Eugene: we plan to remove the tagger system on Friday
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Yi">Yi</span></h1>
            <ul>
              <li> solenoid
                <ul>
                  <li> power supply
                    <ul>
                      <li> 2 spare controller boards came in
                      </li>
                      <li> 1 repaired
                      </li>
                      <li> makes 4 total
                      </li>
                    </ul>
                  </li>
                  <li> slew rate slower than that set
                    <ul>
                      <li> suspicion: either connect of board or fault
                        in daughter board
                      </li>
                      <li> will order a spare daughter board
                      </li>
                      <li> can do a test at low current
                      </li>
                      <li> should not impede operations
                      </li>
                    </ul>
                  </li>
                  <li> Eugene: helium in magnet for a month now
                  </li>
                  <li> Jonathan Creel adjusted parameters on JT valves,
                    smoother operation now
                  </li>
                  <li> Eugene: flow controller for leads on the way
                  </li>
                </ul>
              </li>
              <li> active collimator
                <ul>
                  <li> high-speed EPICS readout implemented
                  </li>
                  <li> analyzer can read from ROOT file
                  </li>
                </ul>
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Mark_I.">Mark I.</span></h1>
            <ul>
              <li> set-up and ran farm jobs for simulation of
                commissioning
                <ul>
                  <li> in support of work by Simon, Paul, Sean
                  </li>
                </ul>
              </li>
              <li> Eugene: Gagik Gavalian has developed an EVIO-to-ROOT
                converter for Hall B
                <ul>
                  <li> we will follow-up to see if we can use it
                  </li>
                </ul>
              </li>
              <li> start counter
                <ul>
                  <li> need TDC readout
                    <ul>
                      <li> TDC read-out not a priority
                      </li>
                      <li> yet TDC is the principle read-out method for
                        this detector
                      </li>
                    </ul>
                  </li>
                  <li> need self-trigger
                    <ul>
                      <li> start counter trigger not a priority
                      </li>
                      <li> yet start counter trigger useful for cosmic
                        testing of many other detector subsystem e. g.,
                        CDC, BCAL (radial track path enhancement), FDC
                        (in coincidence with TOF to give horizontal
                        cosmics) and may be useful for beam data as well
                      </li>
                    </ul>
                  </li>
                </ul>
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Simon">Simon</span></h1>
            <ul>
              <li> continuing work on commissioning geometry
              </li>
              <li> CDC alignment
                <ul>
                  <li> parameter cut-off necessary in fits
                  </li>
                  <li> more data with BCAL as trigger needed
                    <ul>
                      <li> last large run did not have CDC times,
                        neither from TDCs nor from FADCs
                      </li>
                    </ul>
                  </li>
                </ul>
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Mark_D.">Mark D.</span></h1>
            <ul>
              <li> working on BCAL data analysis, low-level
                reconstruction code
                <ul>
                  <li> Eugene: what is z-resolution?
                    <ul>
                      <li> with LED data, using ADC timing, about 3 cm
                      </li>
                    </ul>
                  </li>
                </ul>
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Beni">Beni</span></h1>
            <ul>
              <li> F125 programming
                <ul>
                  <li> will average over 16 samples to determine
                    pedestal
                  </li>
                  <li> one ROC has the new firmware
                  </li>
                </ul>
              </li>
              <li> why flashes cause DAQ to crash?
                <ul>
                  <li> sometimes one ROC does not see trigger
                  </li>
                  <li> need to maintain sync in those cases
                  </li>
                  <li> DMA buffer read-out error only problem on
                    ROC-FDC-10
                  </li>
                </ul>
              </li>
              <li> working on CDC monitoring histograms
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Alexandre">Alexandre</span></h1>
            <ul>
              <li> implement the run plan
                <ul>
                  <li> procedures need work
                  </li>
                </ul>
              </li>
              <li> OSLs being placed in tagger hall[?]
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Elton">Elton</span></h1>
            <ul>
              <li> implementing BCAL config for LED pulsing through
                trigger supervisor, getting help from Beni
              </li>
              <li> cell-phone-carrying personnel cross-training on Wed.
              </li>
              <li> safety walk-through Thu. at 9:30 am
              </li>
              <li> Eugene has a cell phone charging station, Yi will
                take it off his hands and install it in the proper
                location
              </li>
            </ul>
            <h1> <span class="mw-headline" id="Hovanes">Hovanes</span></h1>
            <ul>
              <li> beam profiler is back in position after survey
                <ul>
                  <li> it is downstream of FCAL, and connected, no
                    counts yet
                  </li>
                </ul>
              </li>
              <li> halo counters installed, need cabling to electronics
              </li>
              <li> working on overview gui for EPICS, summary of
                important EPICS channel all in one place
              </li>
              <li> installing some PDU for devices that may need remote
                power cycling
              </li>
              <li> moving test set-up for amorphous radiator harp from
                EEL to tagger hall in case motion is needed for the run
              </li>
              <li> getting access to Accelerator EPICS GUIs
              </li>
              <li> Fernando asked about keys for cabinets, Tom had a
                plan for keybox, Hovanes will check on status
              </li>
            </ul>
          </div>
        </div>
      </div>
    </div>
    <br>
    <pre class="moz-signature" cols="72">-- 
Mark M. Ito, Jefferson Lab, <a class="moz-txt-link-abbreviated" href="mailto:marki@jlab.org">marki@jlab.org</a>, (757)269-5295
</pre>
  </body>
</html>